3. An ideal gas with a constant number of molecules is maintained at a constant pressure. At 370PC, the volume of the gas is120 m? What is the volume of the gas when the temperature is increased to 134 12 marks
Answer #1

Given: T1= 37°c ; T2= 134 °c ; V1= 1.20 m3

Solution:

According to ideal gas law : PV = nRT = NkT

We have,

V2/V1 = T2/T1

V2/1.20= (134/37)

V2 = 4.35 m3